OpenStudy (anonymous):

Combine the factors of (-3x^2y^5) (2xy^3) into the form Cx^ay^b . show work as much as possible, explaing as well. ALOHA:]

OpenStudy (radar):

\[(-3x ^{2}y ^{5})\times(2xy ^{3})=-6x ^{3}y ^{8}\]

OpenStudy (radar):

Have done work on the radar at Mt. Kala and at the indicator site at Diamond Head.

OpenStudy (anonymous):

thank you, but how did you get your answer?

OpenStudy (radar):

It is a multiplication problem. Do it in parts. -3 X 2 =-6,

OpenStudy (radar):

x^2 X x = x^3 (addition of exponents) y^5 X y^3 = y^8 combine parts and get -6x^3y^8